OpenStudy (anonymous):

A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a black marble?

OpenStudy (anonymous):

The probability that Events A and B both occur is the probability of the intersection of A and B. The probability of the intersection of Events A and B is denoted by P(A ∩ B). I'm thinking this is what you need.P(A and B)=P(A)xP(B)... so we get P(blue marble)xP(black marble). P(blue marble)=8/18=4/9. P(black marble)=3/18=1/6. req.probability=4/9x1/6 =4/36=1/9.